Two charged particles having masses in the ratio 2:3 and charges in the ratio 1:2 are released from rest in a uniform electric field.After a time 1 minute their K.E will be in the ratio____________?

Grade:12th pass

1 Answers

Eshan
5 years ago
Force acting on particles=F=qE
HenceF\propto q
\implies a\propto \dfrac{q}{m}
Now,v\propto a
Hencev\propto \dfrac{q}{m}

Therefore, kinetic energyKE\propto m(\dfrac{q^2}{m^2})
KE\propto \dfrac{q^2}{m}
\dfrac{KE_1}{KE_2}=\dfrac{1/4}{2/3}
=\dfrac{3}{8}